# FIVB ranking: Misstep in the right direction

### v1

@@ 3.2 Finding thresholds $c$ and HFA parameter $\eta$
    $\hat{\boldsymbol{c}}=\mathop{\mathrm{argmin}}_{\boldsymbol{c},\eta=0}U(\boldsymbol{c}).$ (59)
@@
    $\hat{\boldsymbol{c}},\hat{\eta}=\mathop{\mathrm{argmin}}_{\boldsymbol{c},\eta}U(\boldsymbol{c},\eta),$ (61)
@@
Figure 2: Results of optimization in (61): thresholds $\hat{c}_{0}$, $\hat{c}_{1}$ and the HFA $\hat{\eta}$. The horizontal dotted lines are drawn for the values $c^{{\textnormal{FIVB}}}_{0}=-1.06$ and $c^{{\textnormal{FIVB}}}_{1}=-0.394$ used by the FIVB ranking, see (23).
@@
  • The FIVB thresholds $\boldsymbol{c}^{{\textnormal{FIVB}}}$ are close, but not identical to, those obtained via optimization. This discrepancy should be attributed to the difference in the data sets from which the thresholds were inferred, but also due to the procedure described by the FIVB ranking, which defined $\boldsymbol{c}$ from the games of the “teams with similar skills”. [^8] We note that this approach to find the model is rather ambiguous as, to find which skills are similar, we have to estimate them, and this requires the model to be defined in the first place.
    In our approach, we used all the games, which may explain some improvement in the average validation metric $U(\boldsymbol{p})$.
  • The decomposition into the games played on the neutral and home venues indicates that, by including the HFA into the model via $\eta>0$, we can improve the prediction for the home games, and, rather unsurprisingly, this modification has practically no impact on the neutral-venue games.
  • Attention should be paid to the metric $V(\boldsymbol{p})$ shown in the right axis, where we see that the improvements are at the order of fraction of a percent. For example, the most significant improvement appears for home-games, where, incorporation of $\eta>0$ and optimization of the thresholds, improve the results, by $\sim 1\%$; i.e., from $V\approx 25.8\%$ to $V\approx 26.8\%$. This can be compared to the performance of the empirical entropy shown in Table 4, which is notably worse.
@@ 3.3.1 From thresholds to numerical scores
which, from $g^{{\textnormal{FIVB}}}_{y}(z_{{\textnormal{o}}};\boldsymbol{r})=\check{r}(z_{{\textnormal{o}}})-r_{y}=-r_{y}$, [^9] Because, due to symmetry, the expected numerical score is zero, $\check{r}(0)=0$ yields
@@
where we used (32) and (63); the notation $\tilde{r}_{y}(\boldsymbol{c})$ emphasizes the dependence of $\tilde{\boldsymbol{r}}$ on the thresholds $\boldsymbol{c}$ which define the CL model. Note also that $\tilde{r}_{y}$ is not uniquely defined, as $\tilde{r}_{0}$ can be arbitrarily fixed, and, for comparison with the FIVB ranking, we set $\tilde{r}_{0}\equiv r^{{\textnormal{FIVB}}}_{0}=2.0$.
Using $\boldsymbol{c}=\boldsymbol{c}^{{\textnormal{FIVB}}}$ in (65) yields
@@
These values are rather different from those shown in Table 2 and, more importantly, when we use them to calculate the implicit loss functions $\ell^{{\textnormal{FIVB}}}_{y}(z;\tilde{\boldsymbol{r}}),y=0,1,2$ (shown, scaled with $\tilde{a}$ and shifted with $\tilde{b}$, in Fig. 4 with dashed-dotted lines), the latter are indistinguishable from $\ell_{y}(z)$ in the vicinity of $z_{{\textnormal{o}}}=0$.
@@ 3.3.2 Numerical optimization
Instead of the analytical approach, shown in the previous section, the numerical optimization of $\boldsymbol{r}$, takes into account the actual outcomes of the games.
@@
    We use $\boldsymbol{c}^{\textnormal{FIVB}}$ and $\boldsymbol{r}^{\textnormal{FIVB}}$ specified by the FIVB ranking, and given, respectively, in (23) and Table 2, and we set $\eta=0$. This is the reference, currently used by the FIVB.
@@
This may appear surprising and counterintuitive, but only if we interpret the numerical score as related to the order of the outcomes. We should remember that ordinal variables do not have intrinsic numerical values, and numerical scores are parameters that allow us to adjust the form of the implicit loss function $\ell^{\text{FIVB}}_{y}(z)$. With such a perspective, the non-monotonic behavior of $r_{y}$ is allowed.111111 We still want to know, if, for $\hat{\boldsymbol{r}}$ given in (69) (where $\hat{r}_{3}>\hat{r}_{2}$) the implicit loss functions $\ell^{{\textnormal{FIVB}}}_{y}(z;\hat{\boldsymbol{r}})$ remain convex in $z$. For this, it suffices to calculate $\check{r}(z)$ and verify that it is monotonically increasing in $z$. In fact, this is the case for all $\tilde{\boldsymbol{r}}$ we obtained. Note that this does not contradict Lemma 1 because the monotonic behavior of $r_{y}$ was a sufficient (but not necessary) condition to ensure the convexity of $\ell^{{\textnormal{FIVB}}}_{y}(z;\boldsymbol{r})$.
@@
Thus, it appears that the numerical scores $\boldsymbol{r}^{{\textnormal{FIVB}}}$ were not formally optimized — a conjecture supported by the fact that their origin is not explained in FIVB (2024).
We speculate that the numerical scores $\boldsymbol{r}$ are defined to be similar to the points in the volleyball leagues, where winning “$3-0$” grants $3$ points to the winner, winning “$3-1$” grants $2$ points, and winning “$3-2$” grants $1$ point. Multiplying these points by $2/3$, this scoring rule would be equivalent to granting $2.0$, $1.33$, and $0.66$ points – values somewhat similar to $r^{{\textnormal{FIVB}}}_{0}=2.0$, $r^{{\textnormal{FIVB}}}_{1}=1.5$, and $r^{{\textnormal{FIVB}}}_{2}=1.0$.
Regardless of the origin of $\boldsymbol{r}^{{\textnormal{FIVB}}}$, it is more sound to see the numerical score $\boldsymbol{r}$ as free parameters which allow us to make the implicit loss function $\ell^{{\textnormal{FIVB}}}_{y}(z;\boldsymbol{r})$ “behave” similarly to the optimal logarithmic loss $\ell_{y}(z)$.
Figure 7: Validation metrics $U^{{\textnormal{ntr}}}$ (hollow markers) and $U^{{\textnormal{hfa}}}$ (colored markers) given by (57)-(58) shown as a function of $\gamma$ using the loss function $\ell_{y}(z)$ defined in (19) and different strategies of fixing the weights $\xi_{v}$ which depend on the games’ categories, including equal weights $\xi_{v}\equiv 1$, $\xi^{{\textnormal{FIVB}}}_{v}$ specified in Table 1, and $\hat{\xi}_{v}$, optimized via (71).
@@ 3.4 Weights
4 Alternative model: CL model with logistic distribution
We conjectured in Sec. 2.2, that the main reason for not directly using the derivative of the log-score is due to the cumbersome formula (32), which not only lacks simple interpretation, but is numerically unstable.
We can now explore an alternative CL model which is simple to implement and also yields an easy-to-interpret SG algorithm. The model is obtained by replacing the Gaussian CDF with the logistic one, i.e., (1) is defined as
$\mathsf{P}^{\mathcal{L}}_{y}(z)=\mathcal{L}(z+c_{y})-\mathcal{L}(z+c_{y-1}),$ (72)
where $\mathcal{L}(z)$ is defined in (8), and the log-score is calculated plugging (72) in (6)
$\ell^{\mathcal{L}}_{y}(z)$ $=-\log\mathsf{P}^{\mathcal{L}}_{y}(z)$ (73)
$=z-\log(\mathrm{e}^{-c_{y-1}}-\mathrm{e}^{-c_{y}})-\log\mathcal{L}(z+c_{y})-\log\mathcal{L}(z+c_{y-1}).$ (74)
Before passing to the numerical optimization, we want to point out the fact that, to find the thresholds $\boldsymbol{c}$, we may reuse the FIVB thresholds $\boldsymbol{c}^{{\textnormal{FIVB}}}$, by exploiting the similarity of the CDF s obtained via scaling
$\Phi(z)\approx\mathcal{L}(z/\sigma),$ (75)
where $\sigma$ is an appropriately chosen scaling. For example, by equating the derivatives of both CDF s at $z_{\textnormal{o}}=0$, i.e.,
$\Phi^{\prime}(z)|_{z=z_{{\textnormal{o}}}}$ $=\frac{1}{\sigma}\mathcal{L}^{\prime}(z/\sigma)|_{z=z_{{\textnormal{o}}}},$ (76)
yields
$\sigma$ $=\sqrt{\pi/8}\approx 0.63.$ (77)
The approximate equivalence of the models (21) and (72), is obtained with the following substitution
$\Phi\big{(}z+\eta+c\big{)}$ $\approx\mathcal{L}\big{(}z/\sigma+\eta/\sigma+c/\sigma\big{)},$ (78)
that is, if $\boldsymbol{c}$ and $\eta$ are the parameters of the model (21), then the logistic CL model (72) should use $\boldsymbol{c}/\sigma$ and $\eta/\sigma$, and the scale $s=\sigma$, see comments on the scale in Sec. 2.
To appreciate the similarity of the models, Fig. 8 shows the functions $\mathsf{P}_{y}(z)$ and $\mathsf{P}^{\mathcal{L}}_{y}(z/\sigma)$ defined respectively in (21) and in (72). Note that, for $\mathsf{P}_{y}(z)$, we use the FIVB parameters $\boldsymbol{c}^{{\textnormal{FIVB}}}$, while in $\mathsf{P}^{\mathcal{L}}_{y}(z/\sigma)$, we use their scaled version $\boldsymbol{c}^{{\textnormal{FIVB}}}/\sigma$.
Figure 8: Models $\mathsf{P}_{y}(z;\boldsymbol{c})$ and $\mathsf{P}^{\mathcal{L}}_{y}(z/\sigma;\boldsymbol{c})$ defined respectively in (21) and (72), obtained from the Gaussian and the logistic CDF s; note the scaling by $\sigma$ applied in the latter.
To assess the CL model (72), we proceed as before, via optimization:
$\hat{\boldsymbol{c}}^{\mathcal{L}},\hat{\eta}^{\mathcal{L}}=\mathop{\mathrm{argmin}}_{\boldsymbol{c},\eta}U(\boldsymbol{c},\eta,\boldsymbol{p}_{\backslash\{\boldsymbol{c},\eta\}}),$ (79)
where the superscript $(\cdot)^{\mathcal{L}}$ is used to denote the parameters used in (72).
The ALO metrics are shown in Fig. 9, where the legend distinguishes the models (21) and (72) by indicating the CDF they use, i.e., $\Phi$ or $\mathcal{L}$. The results obtained using Gaussian CDF (i.e., $\Phi$) are reproduced from Fig. 3.
The main observations are the following:
  • The CL logistic model, with the scaled version of the FIVB thresholds, $\boldsymbol{c}^{{\textnormal{FIVB}}}/\sigma$, outperforms the FIVB model. This is a notable result, obtained without any numerical optimization of the model parameters.
  • Moreover, this choice of parameters is near-optimal. In fact, both the Gaussian and logistic CL models, yield similar performance after optimization.
The above observations indicate that, compared to the FIVB model, the CL logistic model is less sensitive to the imperfect parameter estimation which may be explained by heavier distribution tails of the logistic model; see Fig. 8.
Figure 9: Validation metrics $U^{{\textnormal{ntr}}}$ (hollow markers) and $U^{{\textnormal{hfa}}}$ (colored markers) given by (57)-(58), shown as a function of $\gamma$ for the logistic CL model, (72) (and indicated with $\mathcal{L}$) are compared with the metrics obtained using the Gaussian CL model, (21) (and indicated with $\Phi$).
4.1 SG algorithm
The SG algorithm is obtained by finding the derivative of (73)
$g^{\mathcal{L}}_{y}(z)$ =˙$=\dot{\ell}^{\mathcal{L}}_{y}(z)=\mathcal{L}(z+c_{y})+\mathcal{L}(z+c_{y-1})-1$ (80)
$=\mathcal{L}(z+c_{y-1})-\mathcal{L}(-z-c_{y}).$ (81)
Since the implementation of (81) is straightforward, the SG rating algorithm
$\hat{\boldsymbol{\theta}}_{t+1}$ $\leftarrow\hat{\boldsymbol{\theta}}_{t}-\mu\boldsymbol{x}_{t}g^{\mathcal{L}}_{y_{t}}(\boldsymbol{x}_{t}^{\mathsf{T}}\hat{\boldsymbol{\theta}}_{t})$ (82)
may be used directly without any numerical issues, and we note that it does not make any reference to the numerical scores, i.e., we do not need to assign any value to the observed outputs. This is exactly what we repeatedly said in this paper: ordinal results $y_{t}$, indeed, do not have any intrinsic value.
The interpretation of (82) is quite interesting. Namely, from (24), we know that $\Pr\left\{Y\leq y|z\right\}=\mathcal{L}(z+c_{y})$, and thus,
$\mathcal{L}(z+c_{y-1})=\Pr\left\{Y<y|z\right\}$ (83)
$1-\mathcal{L}(z+c_{y})=\Pr\left\{Y>y|z\right\}$ (84)
so, (81) becomes
$g^{\mathcal{L}}_{y}(z)$ $=\Pr\left\{Y<y|z\right\}-\Pr\left\{Y>y|z\right\}$ (85)
$=\Pr\left\{Y\leq y|z\right\}-\Pr\left\{Y\geq y|z\right\}.$ (86)
Therefore, the algorithm (82) is driven by the difference between the right- and left-hand side distribution tails. Note that, the derivative $g^{\mathcal{L}}_{y_{t}}(z_{t})$ becomes null if the conditional distribution $\Pr\left\{Y_{t}|z_{t}\right\}$ has equal tails, or, equivalently, if $y_{t}$ is a perfect median of $Y_{t}$ (conditioned on $z_{t}$). [^12] For a discrete random variable $Y$, $y$ is its median if $\Pr\left\{Y\leq y\right\}\geq 0.5$ and $\Pr\left\{Y\geq y\right\}\geq 0.5$. We call $y$ a perfect median if $\Pr\left\{Y<y\right\}=\Pr\left\{Y>y\right\}$.
@@ 5 Real-time ranking
We will thus keep the FIVB model defined by the threshold parameters $\boldsymbol{c}^{{\textnormal{FIVB}}}$ and evaluate i) the choice of the numerical score $\boldsymbol{r}$ used in the implicit loss function $\ell^{{\textnormal{FIVB}}}$, and specified in (66), ii) the values of the weights $\boldsymbol{\xi}$, see Sec. 3.4, and iii) the choice of the threshold parameters in the logistic CL model, see Sec. 4. Simply put, we do not carry out any explicit optimization of the model parameters when using the real-time ranking but, rather, rely on the parameters obtained from analysis. In this way, we avoid the contentious issue of choosing the model parameters from the data. The only exception is the choice of the HFA parameter which we set as $\eta=0.2$.
To initialize the SG algorithm (15), we use the skills $\theta_{m,0},m=1,\ldots,M$, where $\theta_{m,0}$ is read from the official FIVB ranking of the team $m$ at the time of their first game after 2020.131313Thus, we do not take into account the fact that the FIVB ranking penalizes “inactive” teams, i.e., those that do not play any games in a given year, and whose rating is then reduced by 50 points. By initializing the skills with those provided by the official ranking allows us to deal with the practical aspect of switching from one ranking (here, the official one) to another (the one we propose).
In order to switch the model from the Gaussian to the logistic CDF, we need to change the scaling, but this is also handled analytically: to obtain a quasi-equivalence of the model, the logistic model (72) needs to employ scaling $s^{\mathcal{L}}=\sigma s\approx 80$ (where $s=125$, see (21)) and use $\eta^{\mathcal{L}}=s/\sigma\approx 0.12$. Note that a similar strategy for adjusting the scales to avoid convergence problems is also discussed in (Szczecinski and Roatis, 2022, Sec. 5.1), but it is based on numerical results, while we do it analytically.
@@
These metrics are, in essence, equivalent to those shown in (57)-(58), which we have shown in the figures. The difference is that now skills $\hat{\boldsymbol{\theta}}_{t}$ are estimated using the SG algorithm.
@@
$\ell^{{\textnormal{FIVB}}}$ $\boldsymbol{c}^{{\textnormal{FIVB}}},\boldsymbol{r}^{{\textnormal{FIVB}}},\boldsymbol{\xi}^{{\textnormal{FIVB}}},\eta=0$ 0.01 1.52 1.51 1.53 0.94
@@
$\ell^{{\textnormal{FIVB}}}$ $\boldsymbol{c}^{{\textnormal{FIVB}}},\boldsymbol{r}^{{\textnormal{FIVB}}},\boldsymbol{\xi}^{{\textnormal{FIVB}}},\eta=0.2$ 0.01 1.52 1.51 1.52 0.94
@@
$\ell^{{\textnormal{FIVB}}}$ $\boldsymbol{c}^{{\textnormal{FIVB}}},\tilde{\boldsymbol{r}},\boldsymbol{\xi}^{{\textnormal{FIVB}}},\eta=0.2$ 0.01 1.52 1.51 1.53 0.94
@@
$\ell^{{\textnormal{FIVB}}}$ $\boldsymbol{c}^{{\textnormal{FIVB}}},\boldsymbol{r}^{{\textnormal{FIVB}}},\xi_{v}=1,\eta=0.2$ 0.10 1.48 1.49 1.45 0.87
$\ell^{{\textnormal{FIVB}}}$ $\boldsymbol{c}^{{\textnormal{FIVB}}},\tilde{\boldsymbol{r}},\xi_{v}=1,\eta=0.2$ 0.10 1.47 1.48 1.44 0.88
$\ell$ $\boldsymbol{c}^{{\textnormal{FIVB}}},\xi_{v}=1,\eta=0.2$ 0.20 1.46 1.48 1.43 0.85
$\ell^{\mathcal{L}}$ $\boldsymbol{c}^{{\textnormal{FIVB}}}/\sigma,\xi_{v}=1,\eta=0.12,s=80$ 0.40 1.45 1.46 1.42 0.88
Table 5: The metrics (87)-(89) obtained using the SG algorithm using different loss functions and parameters. We always show the results with the step $\hat{\mu}$ obtained via (90), except in the first three cases where we also show the results obtained using $\mu=0.01$ which is defined in the FIVB ranking.
@@
Note that even using exactly the same parameters as the FIVB ranking (the first line in Table 5), our ranking is not the same as the official one because we discarded the forfeited games.
The results validate the observations we made analyzing the models. Namely, i) an improvement in performance is obtained using the parameter HFA $\eta=0.2$ and the numerical scores $\tilde{r}_{y}(\boldsymbol{c}^{{\textnormal{FIVB}}})$, see (65), and ii) the weighting with $\xi_{v}>1$ is counterproductive.

### v2

@@ 3.2 Finding thresholds $c$ and HFA parameter $\eta$
[…]
@@
[…]
@@
Figure 2: Results of optimization in (42): thresholds $\hat{c}_{0}$, $\hat{c}_{1}$ and the HFA $\hat{\eta}$. The horizontal dotted lines are drawn for the values $c^{{\textnormal{FIVB}}}_{0}=-1.06$ and $c^{{\textnormal{FIVB}}}_{1}=-0.394$ used by the FIVB ranking, see (18). All thresholds satisfy (17) so $c_{2}\equiv 0$ need not be shown.
@@
From Fig. 2, we see that the FIVB thresholds $\boldsymbol{c}^{{\textnormal{FIVB}}}$ are close to, but not identical with those obtained through optimization. This discrepancy should be attributed to the difference in the data sets from which the thresholds were inferred, but also to the procedure described by the FIVB ranking, which defined $\boldsymbol{c}^{{\textnormal{FIVB}}}$ from the matches of the “teams with similar skills”. [^10] We note that this approach to find the model is rather ambiguous as, to find which skills are similar, we have to estimate them, and this requires the model to be defined in the first place. In our approach, we used all matches, which may explain some improvement in the validation metrics $U(\hat{\boldsymbol{p}})$ we observe in Fig. 3.
From Fig. 3 we observe that:
  • The decomposition into the matches played on the neutral and home venues indicates that, by including the HFA into the model via $\eta$, we can improve the prediction for the home matches, and, rather unsurprisingly, this modification has practically no impact on the neutral-venue matches.
  • Attention should be paid to the metric $V(\boldsymbol{p})$ shown on the right axis, where we see that the improvements are on the order of a fraction of a percent. For example, the most significant improvement appears for home matches, where, by optimizing $\eta$ and the thresholds $\boldsymbol{c}$ in (42), the results are improved by $\sim 1\%$; i.e., from $V\approx 25.8\%$ to $V\approx 26.8\%$.
@@ 3.3.1 From thresholds to numerical scores
which, from $g^{{\textnormal{FIVB}}}_{y}(z_{{\textnormal{o}}};\boldsymbol{r})=\check{r}(z_{{\textnormal{o}}})-r_{y}=-r_{y}$, [^11] Because, due to symmetry, the expected numerical score is zero, i.e., $\check{r}(0)=0$ yields
@@
where we used (27) and (44); the notation $\tilde{r}_{y}(\boldsymbol{c})$ emphasizes the dependence of $\tilde{\boldsymbol{r}}$ on the thresholds $\boldsymbol{c}$ which define the CL model so (46) is valid for any $\boldsymbol{c}$. Note also that $\tilde{r}_{y}$ is not uniquely defined because we can arbitrarily fix $\tilde{r}_{0}$, and, for comparison with the FIVB ranking, we set $\tilde{r}_{0}\equiv r^{{\textnormal{FIVB}}}_{0}=2.0$.
We can now calculate the scores by applying (46) to the FIVB-defined thresholds, $\boldsymbol{c}=\boldsymbol{c}^{{\textnormal{FIVB}}}$, which yields $\tilde{\boldsymbol{r}}=[\tilde{r}_{0}(\boldsymbol{c}^{{\textnormal{FIVB}}}),\ldots,\tilde{r}_{L-1}(\boldsymbol{c}^{{\textnormal{FIVB}}})]$
@@
These values are rather different from those shown in Table 2 and, more importantly, when we use them to calculate the implicit loss functions $\ell^{{\textnormal{FIVB}}}_{y}(z;\tilde{\boldsymbol{r}}),y=0,1,2$ (shown, scaled with $\tilde{a}$ and shifted with $\tilde{b}_{y}$, in Fig. 4 with dashed-dotted lines), the latter are indistinguishable from $\ell_{y}(z)$ in the vicinity of $z_{{\textnormal{o}}}=0$. Clearly, when compared to the implicit loss $\ell_{y}^{{\textnormal{FIVB}}}(z;\boldsymbol{r}^{{\textnormal{FIVB}}})$ with FIVB-defined scores $\boldsymbol{r}^{{\textnormal{FIVB}}}$, the loss $\ell^{\textnormal{FIVB}}_{y}(z;\tilde{\boldsymbol{r}})$ offers a better approximation of the log-loss $\ell_{y}(z)$. And this improvement is obtained solely using the numerical score $\tilde{\boldsymbol{r}}$.
@@ 3.3.2 Numerical optimization
Instead of the analytical approach, shown in the previous section, the numerical optimization of $\boldsymbol{r}$, takes into account the actual outcomes of the matches.
@@
    We use $\boldsymbol{c}^{\textnormal{FIVB}}$ and $\boldsymbol{r}^{\textnormal{FIVB}}$ specified by the FIVB ranking, and given, respectively, in (18) and Table 2, and we set $\eta=0$. This is the reference, currently used by FIVB.
@@
This may appear surprising and counterintuitive, but only if we interpret the numerical score as related to the order of the outcomes. We should remember that ordinal variables do not have intrinsic numerical values, and numerical scores are parameters that allow us to adjust the form of the implicit loss function $\ell^{\text{FIVB}}_{y}(z)$. With such a perspective, the non-monotonic behavior of $r_{y}$ is allowed.131313 We still want to know, if, for $\hat{\boldsymbol{r}}$ given in (51) (where $\hat{r}_{3}>\hat{r}_{2}$) the implicit loss functions $\ell^{{\textnormal{FIVB}}}_{y}(z;\hat{\boldsymbol{r}})$ remain convex in $z$. For this, it suffices to calculate $\check{r}(z)$ and verify that it is monotonically increasing in $z$. In fact, this is the case for all $\hat{\boldsymbol{r}}$ we obtained. Note that this does not contradict Lemma 1 because the monotonic behavior of $r_{y}$ was a sufficient (but not necessary) condition to ensure the convexity of $\ell^{{\textnormal{FIVB}}}_{y}(z;\boldsymbol{r})$.
@@
Thus, it appears that the numerical scores $\boldsymbol{r}^{{\textnormal{FIVB}}}$ were not formally optimized — a conjecture supported by the fact that their origin is not explained in (FIVB, 2024). However, regardless of the origin of $\boldsymbol{r}^{{\textnormal{FIVB}}}$, it is more sound to see the numerical score $\boldsymbol{r}$ as free parameters which allow us to make the implicit loss function $\ell^{{\textnormal{FIVB}}}_{y}(z;\boldsymbol{r})$ “behave” similarly to the optimal logarithmic loss $\ell_{y}(z)$.
Figure 7: Validation metrics $U^{{\textnormal{ntr}}}$ (hollow markers) and $U^{{\textnormal{hfa}}}$ (colored markers) given by (36)-(37) shown as a function of $\gamma$ using the loss function $\ell_{y}(z)$ defined in (14) and different strategies of fixing the weights $\xi_{v}$ which depend on the matches’ categories, including equal weights $\xi_{v}\equiv 1$, $\xi^{{\textnormal{FIVB}}}_{v}$ specified in Table 1, and $\hat{\xi}_{v}$, optimized via (52).
@@ 3.4 Weights
To clarify the somewhat intriguing behavior of $U(\boldsymbol{p})$ for the optimized results (marked with stars in Fig. 7), which, for large $\gamma$, is notably better than in the case of $\xi_{v}\equiv 1$, we write the optimization (35) as
$\hat{\boldsymbol{\theta}}_{\backslash{t}}$ $=\mathop{\mathrm{argmin}}_{\boldsymbol{\theta}}\Big{[}\sum_{\begin{subarray}{c}\tau=1\\ \tau\neq t\end{subarray}}^{T}\hat{\xi}_{v_{\tau}}(\hat{\gamma})\ell^{{\textnormal{loss}}}_{y_{\tau}}(\boldsymbol{x}_{\tau}^{\mathsf{T}}\boldsymbol{\theta})+\hat{\gamma}\|\boldsymbol{\theta}\|^{2}\Big{]}.$ (53)
where $\hat{\xi}_{v}(\hat{\gamma})$ are the optimal weights obtained in (52) for an optimal $\hat{\gamma}\approx 0.5$. Since the multiplication of the cost function by $\alpha>0$ is irrelevant to the optimization results, using weights $\alpha\hat{\xi}_{v}(\gamma)$ and the regularization parameter $\gamma=\alpha\hat{\gamma}$, will not change $\hat{\boldsymbol{\theta}}_{\backslash{t}}$. In other words, increasing $\gamma$ we can simultaneously increase $\boldsymbol{\xi}$ and maintain performance $U(\boldsymbol{p})$ flat in $\gamma$. The only reason it does not happen is because we impose the constraint $\xi_{0}=1$.
@@ 4 Real-time ranking
We will thus keep the FIVB model defined by the threshold parameters $\boldsymbol{c}^{{\textnormal{FIVB}}}$ and evaluate i) the choice of the numerical score $\boldsymbol{r}$ used in the implicit loss function $\ell^{{\textnormal{FIVB}}}$, and specified in (47) and ii) the values of the weights $\boldsymbol{\xi}$, see Sec. 3.4, Simply put, we do not carry out any explicit optimization of the model parameters when using the real-time ranking but, rather, rely on the parameters obtained from analysis. In this way, we avoid the contentious issue of choosing the model parameters from the data. The only exception is the choice of the HFA parameter which we set as $\eta=0.2$.
To initialize the SG algorithm (10), we use the skills $\theta_{m,0},m=1,\ldots,M$, where $\theta_{m,0}$ is read from the official FIVB ranking of the team $m$ at the time of their first match after 2020.141414Thus, we do not take into account the fact that the FIVB ranking penalizes “inactive” teams, i.e., those that do not play any matches in a given year, and whose ranking is then reduced by 50 points. By initializing the skills with those provided by the official ranking allows us to deal with the practical aspect of switching from one ranking (here, the official one) to another (the one we propose).
@@
These metrics are, in essence, equivalent to those in (36)-(37), which we have shown in the figures. The difference is that now skills $\hat{\boldsymbol{\theta}}_{t}$ are estimated using the SG algorithm.
@@
A $\ell^{{\textnormal{FIVB}}}$ $\boldsymbol{c}^{{\textnormal{FIVB}}},\boldsymbol{r}^{{\textnormal{FIVB}}},\boldsymbol{\xi}^{{\textnormal{FIVB}}},\eta=0$ 0.01 1.52 1.51 1.53 0.94
@@
B $\ell^{{\textnormal{FIVB}}}$ $\boldsymbol{c}^{{\textnormal{FIVB}}},\boldsymbol{r}^{{\textnormal{FIVB}}},\boldsymbol{\xi}^{{\textnormal{FIVB}}},\eta=0.2$ 0.01 1.52 1.51 1.52 0.94
@@
C $\ell^{{\textnormal{FIVB}}}$ $\boldsymbol{c}^{{\textnormal{FIVB}}},\tilde{\boldsymbol{r}},\boldsymbol{\xi}^{{\textnormal{FIVB}}},\eta=0.2$ 0.01 1.52 1.51 1.53 0.94
@@
D $\ell^{{\textnormal{FIVB}}}$ $\boldsymbol{c}^{{\textnormal{FIVB}}},\boldsymbol{r}^{{\textnormal{FIVB}}},\xi_{v}\equiv 1,\eta=0.2$ 0.10 1.48 1.49 1.45 0.87
E $\ell^{{\textnormal{FIVB}}}$ $\boldsymbol{c}^{{\textnormal{FIVB}}},\tilde{\boldsymbol{r}},\xi_{v}\equiv 1,\eta=0.2$ 0.10 1.47 1.48 1.44 0.88
F $\ell$ $\boldsymbol{c}^{{\textnormal{FIVB}}},\xi_{v}\equiv 1,\eta=0.2$ 0.20 1.46 1.48 1.43 0.85
Table 4: The metrics (54)-(56) obtained using the SG algorithm using different loss functions and parameters. We always show the results with the step $\hat{\mu}$ obtained via (57), except in the cases A, B and C, where we also show the results obtained using $\mu=0.01$ which is defined in the FIVB ranking.
@@
Note that even using exactly the same parameters as the FIVB ranking (case A in Table 4), our results are not the same as the official ones because we discarded the forfeited matches; this explains why the Spearman correlation is the largest among the algorithms, but it is not perfect, i.e., $\overline{\rho}<1$.
Without surprise, the best result $\overline{U}=1.46$, is obtained using the true log-score $\ell_{y}(z)$, given by (26) (case F in Table 4) and, on average, this ranking is the least correlated with the FIVB ranking ($\overline{\rho}=0.85$); remember, however, that the implementation of (27) is numerically complex. The second-best result is obtained using i) the numerical scores $\tilde{r}_{y}(\boldsymbol{c}^{{\textnormal{FIVB}}})$, see (46), together with the constant weighting $\xi_{v}\equiv 1$ (case E in Table 4). An improvement in performance from using the HFA $\eta=0.2$ is slight (see cases A and B) but, quantitatively, in line with the improvement observed in home-matches in Fig. 3 (where, after applying the HFA, the value of $U(\boldsymbol{p})$ changes from $1.36$ to $1.34$).
To show an example of how the algorithms (A, E and F) affect the ranking of the top teams, we show the ranking in Table 5, where the three front-runners stay the same, but the position of the remaining teams changes.
case top teams and skills
A ($\mu=0.01$) POL USA JPN BRA ITA ARG RUS
423.8 396.8 345.9 345.0 344.3 317.0 315.7
A ($\mu=0.03$) POL USA JPN ARG SLO GER BRA
517.4 478.2 420.2 374.9 368.1 361.3 360.2
E POL USA JPN ARG ITA SLO GER
468.9 462.2 409.8 381.0 375.3 366.4 356.0
F POL USA JPN GER ARG SLO ITA
528.9 524.3 475.3 436.9 430.6 416.3 401.4
Table 5: Ranking of the top teams in the last day of 2023 for different algorithm with parameters specified in Table 4.